OpenStudy (anonymous):

Solve the system of equations using the substitution method. enter the answer as an ordered pair. 2x-y=8 y=x-3 please help thanks

OpenStudy (anonymous):

substitute y = x-3 in 2x = y -3 for which will help u find x

OpenStudy (anonymous):

okay and how do i find y?

OpenStudy (anonymous):

u find x then substitue x in y = x-3

OpenStudy (anonymous):

ok. got it thanks! :)

OpenStudy (anonymous):

no prbs

OpenStudy (anonymous):

is the correct answer (5, 2) ?

OpenStudy (anonymous):

yup